### 03 Apr 2026 v2

# Achieving double-logarithmic precision dependence in optimization-based quantum unstructured search

Abstract
Grover’s algorithm is a fundamental quantum algorithm that achieves a quadratic speedup for unstructured search problems of size $N$. Recent studies have reformulated this task as a maximization problem on the unitary manifold and solved it via linearly convergent Riemannian gradient ascent (RGA) methods, resulting in a complexity of $\mathcal{O}(\sqrt{N/M}\log(1/\varepsilon))$, where $M$ denotes the number of target items. In this work, we adopt the Riemannian modified Newton (RMN) method to solve the quantum search problem, under the assumption that the ratio $M/N$ is known. We show that, in this setting, the Riemannian Newton direction is collinear with the Riemannian gradient in the sense that the Riemannian gradient is always an eigenvector of the corresponding Riemannian Hessian. As a result, without additional overhead, the proposed RMN method numerically achieves a quadratic convergence rate with respect to the error $\varepsilon$, implying a complexity of $\mathcal{O}(\sqrt{N/M}\log\log(1/\varepsilon))$. Furthermore, our approach remains Grover-compatible, namely, it relies exclusively on the standard Grover diffusion and oracle operators to ensure algorithmic implementability, and its parameter update process can be efficiently precomputed on classical computers.
I Introduction
Unstructured search, the task of finding specific target elements in an unsorted space of size $N$, is a ubiquitous challenge across cryptographic analysis [17, 5], optimization [13, 15], and machine learning [11, 12]. Let $M\ll N$ denote the number of target items. Classically, this problem is inherently inefficient, as guaranteeing success requires $\Omega(N/M)$ queries [4]. This constitutes a severe computational bottleneck for large-scale data retrieval [29, 42].
Grover’s quantum search algorithm resolved this bottleneck by achieving an optimal query complexity of $\Theta(\sqrt{N/M})$ [18, 45, 56, 3], marking a landmark demonstration of quantum advantage. Beyond its direct application to the search problem, Grover’s mechanism is a fundamental primitive in quantum computing. Its principles form the basis of amplitude amplification [7, 49] and have been further generalized within the broader framework of quantum singular value transformation (QSVT) [16, 43].
To gain deeper insights into the mechanics of Grover’s quadratic speedup, researchers have explored various mathematical interpretations. Early efforts characterized Grover’s iteration as a geodesic in complex projective space [44]. Other perspectives have used information geometry to describe the algorithm as a path under the Wigner-Yanase metric [9], while recent work has proposed a connection to imaginary-time evolution (ITE) [50]. These interpretations link the efficiency of quantum search to the geometry of the quantum state space.
A promising modern approach to analyzing such quantum processes is the (Riemannian) manifold optimization framework [1, 23, 6]. Since quantum operations are constrained by unitarity, they naturally lie on the unitary manifold, namely, the unitary group $\{U\in\mathbb{C}^{N\times N}|U^{\dagger}U=I\}$ [20]. Formulating various quantum tasks as optimization problems on these Riemannian manifolds has led to significant progress in quantum tomography [22, 35], ground-state preparation [53, 40, 41, 48, 32], Hamiltonian simulation [30, 34], and circuit compilation [39, 38, 54].
Specifically, a recent study [31] reformulates Grover’s search as a maximization problem on the unitary manifold and proposes a physically implementable Riemannian gradient ascent (RGA) method to achieve the desired quadratic speedup in size $N$. While this first-order approach establishes a rigorous link between the optimization framework and quantum dynamics, it is limited to linear convergence with respect to the accuracy parameter $\varepsilon$, with total complexity $\mathcal{O}(\sqrt{N/M}\log(1/\varepsilon))$.
To further accelerate convergence in terms of the target accuracy $\varepsilon$, second-order optimization methods, such as the Riemannian Newton method [2, 1], provide a natural approach. In this work, we propose a Riemannian modified Newton (RMN) method for Grover’s search problem, where each iteration admits a physically implementable second-order update using only the diffusion and oracle operators, analogous to Grover’s algorithm.
Typically, moving from first-order methods to second-order methods trades faster convergence for higher computational cost per iteration, often due to Hessian inversion. A key theoretical result of our RMN method is that the Riemannian gradient of our cost function in unstructured search is always an eigenvector of the associated Riemannian Hessian. This implies that the Newton direction reduces to a scaled gradient, eliminating the need for explicit Hessian inversion. Consequently, second-order updates incur no additional overhead compared to first-order updates.
We test our RMN method numerically and find that, once the iterates get close to the target state, RMN can attain very high-precision solutions in very few iterations. Furthermore, our numerical results show that RMN achieves a quadratic convergence rate with respect to the accuracy $\varepsilon$, with total complexity $\mathcal{O}(\sqrt{N/M}\log\log(1/\varepsilon))$. However, we emphasize that, in the RMN method, the rotation angle at each iteration must be precomputed on a classical computer under the assumption that the exact value of the ratio $\tau=M/N$ is known. Indeed, when $\tau$ is known, there exist variants of Grover’s algorithm [21, 37, 7] that achieve success probability exactly equal to $1$, with complexity independent of $\varepsilon$. Hence, the primary contribution of this work is not to improve the practical complexity over the existing search algorithms, but rather to provide a new optimization-based method and perspective for quantum unstructured search.
Organization.
In Section˜II, we introduce the preliminaries of the unstructured search problem. In Section˜III, we revisit the Grover-compatible Riemannian gradient ascent method. In Section˜IV, we derive the Riemannian Hessian and present the proposed Grover-compatible Riemannian modified Newton method. Numerical experiments are provided in Section˜V. We conclude in Section˜VI.
II Preliminaries
We first review the notation for the unstructured search problem and then reformulate it as a maximization problem on the unitary manifold, followed by a discussion of the corresponding optimality conditions. In the subsequent sections, we will propose several algorithms.
II.1 Problem statement
Let $\mathcal{H}$ be the $N$-dimensional Hilbert space of an $n$-qubit system with $N=2^{n}$. We consider an unstructured search problem over the finite set $N]:=\{0,1,\ldots,N-1\},$ where a binary oracle function $g\colon[N]\to\{0,1\}$ specifies whether an element is marked. The set of marked items is $S:=\{x\in[N]:g(x)=1\}$ with $M:=|S|$ and typically $1\leq M\ll N$. The goal of the search task is to identify at least one element in $S$.
Let $\{|j\rangle\}_{j=0}^{N-1}$ denote the computational basis of $\mathcal{H}$, where $|j\rangle$ represents the $j$-th standard basis vector. Define the marked subspace as $\mathcal{T}:=\operatorname{span}_{\mathbb{C}}\{\,|x\rangle:x\in S\,\}\subseteq\mathcal{H}$. The quantum search algorithm can then be interpreted as the task of preparing a normalized state vector in $\mathcal{T}$. Let the orthogonal projector onto $\mathcal{T}$ be
$H:=\sum_{x\in S}|x\rangle\langle x|.$ (1)
Grover’s algorithm begins with the uniform superposition over the entire computational basis, i.e., $|\psi_{0}\rangle:=\frac{1}{\sqrt{N}}\sum_{x=0}^{N-1}|x\rangle$. We denote the associated rank-one projector by $\psi_{0}:=|\psi_{0}\rangle\langle\psi_{0}|$. Grover’s algorithm uses two types of quantum gates, the diffusion operator and the oracle operator, defined respectively as
$D(\alpha)$ $:=e^{i\alpha\psi_{0}}=I+\left(e^{i\alpha}-1\right)\psi_{0},$ (2)
$O_{g}(\beta)$ $:=e^{i\beta H}=I+\left(e^{i\beta}-1\right)H,$ (3)
where the parameters $\alpha,\beta\in\mathbb{R}$ denote the angles. Grover iterations are constructed by alternating these two operators, $G(\alpha_{k},\beta_{k})=-D(\alpha_{k})O_{g}(\beta_{k})$. After $T$ iterations, the resulting state is $|\psi_{\text{final}}\rangle=\prod_{k=1}^{T}G(\alpha_{k},\beta_{k})\,|\psi_{0}\rangle$, which aims to approximate the ideal target state
$|\psi^{\star}\rangle:=\frac{1}{\sqrt{M}}\sum_{x\in S}|x\rangle\in\mathcal{T}.$ (4)
The query complexity of Grover’s algorithm is commonly measured by the number of oracle calls $O_{g}(\beta)=e^{i\beta H}$. Detailed circuit implementations of both gates can be found in standard textbooks [27, 45].
In the original version of Grover’s algorithm [18], the parameters are chosen as $\alpha_{k}=\beta_{k}=\pi$ for all $k$. However, this choice may lead to the overshooting problem [8]: performing too many Grover iterations may cause the state to rotate past the target rather than toward it. To mitigate this issue, the $\pi/3$ algorithm [19] fixes the angles at $\alpha_{k}=\beta_{k}=\pi/3$, although this modification loses the quadratic speedup. Later, the fixed-point algorithm [55] was proposed to avoid overshooting while still retaining the quadratic speedup.
II.2 Manifold optimization formulation
In this subsection, we reformulate the quantum search problem as a maximum-eigenstate preparation problem, which can further be cast as a manifold optimization problem. This idea has been proposed in [50, 31]. Throughout, we equip the matrix space with the Frobenius inner product $\langle A,B\rangle:=\operatorname{Tr}(A^{\dagger}B)$.
Let us consider the Hamiltonian $H$ defined in (1). It satisfies $H=H^{\dagger}=H^{2}$, which implies that its eigenvalues can only take the values 0 and 1. In fact, the spectrum of $H$ consists of eigenvalue 1 on the marked subspace $\mathcal{T}$ with multiplicity $M$, and eigenvalue 0 on its orthogonal complement $\mathcal{T}^{\perp}$. In particular, we have $H\left|\psi^{\star}\right\rangle=\left|\psi^{\star}\right\rangle.$
Let $q_{0}:=\frac{M}{N}\in(0,1)$ denote the probability of selecting a marked item at random. Applying $H$ to the initial state gives $H\left|\psi_{0}\right\rangle=\sqrt{q_{0}}\left|\psi^{\star}\right\rangle.$ Thus, the initial expectation value satisfies $\left\langle\psi_{0}\right|H\left|\psi_{0}\right\rangle=\|H\left|\psi_{0}\right\rangle\|^{2}=\|\sqrt{q_{0}}\left|\psi^{\star}\right\rangle\|^{2}=q_{0}.$ Indeed, for any state $|\psi\rangle$, the expectation value $\langle\psi|H|\psi\rangle=\sum_{x\in S}|\langle x|\psi\rangle|^{2}\in[0,1]$ represents the probability of observing a marked item when measuring $|\psi\rangle$ in the computational basis.
Based on the above discussion, our goal now is to construct a quantum state whose expectation value with respect to $H$ (i.e., the probability of observing a marked item) equals 1. To achieve this, we start from an easily prepared state $\left|\psi_{0}\right\rangle$ and aim to design a quantum circuit $U$, whose quantum gates are easily implementable on quantum hardware (e.g., diffusion/oracle operators), such that the resulting state $U\left|\psi_{0}\right\rangle$ becomes an eigenstate of $H$ corresponding to its largest eigenvalue 1.
Consequently, we consider the following manifold optimization problem:
$\max_{U\in\mathrm{U}(N)}f(U),$ (5)
where the feasible region is the compact Lie group of $N\times N$ unitary matrices, which forms a Riemannian manifold,
$\mathrm{U}(N):=\left\{U\in\mathbb{C}^{N\times N}\mid U^{\dagger}U=I_{N}\right\},$ (6)
and the cost function $f:\mathrm{U}(N)\rightarrow\mathbb{R}$ is defined by
$f(U):=\operatorname{Tr}\!\left(HU\psi_{0}U^{\dagger}\right)=\langle\psi_{0}|U^{\dagger}HU|\psi_{0}\rangle.$ (7)
To solve the above problem on the manifold $\mathrm{U}(N)$, we will employ geometric tools and algorithms from the manifold optimization framework [1, 6]. In the main text that follows, we only briefly describe the geometric tools on $\mathrm{U}(N)$. A detailed introduction to the geometry of the unitary manifold is provided in Appendix˜A.
To characterize the directions of infinitesimal motion on the manifold $\mathrm{U}(N)$, we introduce the tangent space. For any $U\in\mathrm{U}(N)$, the tangent space at $U$ is given by
$T_{U}:=\left\{\Omega U:\Omega^{\dagger}=-\Omega\right\}=\mathfrak{u}(N)U,$ (8)
where
$\mathfrak{u}(N):=\left\{\Omega\in\mathbb{C}^{N\times N}:\Omega^{\dagger}=-\Omega\right\}$ (9)
denotes the Lie algebra of $\mathrm{U}(N)$, i.e., the real vector space of all $N\times N$ skew-Hermitian matrices. Intuitively, the tangent space can be viewed as the tangent plane to the manifold at a given point (see Fig.˜1).
For an unconstrained optimization problem, it is well known that the optimality condition is that the gradient vanishes. When the problem is restricted to a manifold, this condition becomes the vanishing of the Riemannian gradient (see Appendix˜A for more). Let $\psi_{U}:=U\psi_{0}U^{\dagger}$ denote the density operator corresponding to the output state of the circuit $U$. The Riemannian gradient of $f$ at some $U\in\mathrm{U}(N)$ is given by
$\operatorname{grad}f(U)=[H,\psi_{U}]\,U\in T_{U},$ (10)
where [A,B]:=ABBA[A,B]:=AB-BA. For convenience, we denote the corresponding skew-Hermitian part by
$\widetilde{\operatorname{grad}}f(U)=[H,\psi_{U}]\in\mathfrak{u}(N).$ (11)
When discussing the tangent space $T_{U}=\mathfrak{u}(N)U$, we often focus on $\mathfrak{u}(N)$ itself and omit the right unitary $U$. [31, Theorem 3.2] has shown that $\widetilde{\operatorname{grad}}f\left(U^{\star}\right)=0$ if and only if $U^{\star}\in\mathrm{U}(N)$ globally minimizes $f(U)$ over $\mathrm{U}(N)$ with the minimum value 0, or globally maximizes it with the maximum value 1.
Remark 1.
Note that the above global optimality condition, resembling that of a convex function, is quite special. It holds only for Hamiltonians satisfying $H=H^{2}$. In general cases, it serves merely as a necessary condition for a local extremum.
III Grover-compatible Riemannian gradient ascent method
In this section, we introduce the Grover-compatible Riemannian gradient ascent (RGA) method for problem Eq.˜5. The term gradient ascent indicates that it is a typical first-order optimization algorithm. The term Grover-compatible means that each iteration is implemented using the diffusion and oracle operators (Eqs.˜2 and 3) from Grover’s algorithm, making the method easy to realize on quantum hardware. This RGA method, originally proposed in the author’s previous work [31], serves as a key foundation for the second-order method, the Grover-compatible Riemannian modified Newton method, developed in this work.
III.1 From Euclidean iterations to manifold iterations
Recall that in the classical Euclidean setting, to solve the optimization problem $\max_{x\in\mathbb{R}^{n}}f(x),$ one typically constructs a sequence $\left\{x_{k}\right\}\subseteq\mathbb{R}^{n}$ from an initial iterate $x_{0}\in\mathbb{R}^{n}$ using the update rule
$x_{k+1}=x_{k}+t_{k}\eta_{k},$ (12)
where $\eta_{k}\in\mathbb{R}^{n}$ is a search direction (e.g., usual gradient $\nabla f(x_{k})$) and $t_{k}>0$ is a step size. To extend the update scheme Eq.˜12 to manifolds, the search direction $\eta_{k}$ is now chosen in the tangent space, and a mapping called retraction is used to map $t_{k}\eta_{k}$ back to the manifold.
Figure 1: Schematic illustration of a manifold optimization iteration on the unitary manifold $\mathrm{U}(N)$. Starting from the current point $U_{k}$, a tangent direction $\eta_{k}\in T_{U_{k}}$ (e.g., the Riemannian gradient) is chosen in the tangent space. The retraction $\mathrm{R}_{U_{k}}$ then maps the scaled tangent vector $t_{k}\eta_{k}$ back onto the manifold, producing the next iterate $U_{k+1}=\mathrm{R}_{U_{k}}(t_{k}\eta_{k})$.
Specifically, to solve the manifold optimization problem $\max_{U\in\mathrm{U}(N)}f(U)$, starting from an initial point $U_{0}\in\mathrm{U}(N)$, a sequence $\{U_{k}\}\subseteq\mathrm{U}(N)$ is generated via update rule (see Fig.˜1 for a schematic illustration):
$U_{k+1}=\mathrm{R}_{U_{k}}\!\left(t_{k}\eta_{k}\right),$ (13)
where $\eta_{k}\in T_{U_{k}}$ is a tangent vector at the current point (e.g., Riemannian gradient $\eta_{k}=\operatorname{grad}f(U_{k})$), and $t_{k}>0$ is still a step size. The mapping $\mathrm{R}_{U}:T_{U}\rightarrow\mathrm{U}(N)$ is called a retraction, which enables movement along tangent directions while keeping iterates on the manifold. Formally, a retraction is a smooth mapping $\mathrm{R}_{U}:T_{U}\rightarrow\mathrm{U}(N)$ such that the induced curve $t\mapsto\gamma(t):=\mathrm{R}_{U}(t\eta)$ satisfies
$\gamma(0)=U,\quad\dot{\gamma}(0)=\eta,\quad\forall\eta\in T_{U}.$ (14)
A natural example of a retraction on $\mathrm{U}(N)$ is the Riemannian exponential map, defined by
$\mathrm{R}_{U}^{\text{exp}}(\eta)=e^{\widetilde{\eta}}U,\quad\widetilde{\eta}:=\eta U^{\dagger}\in\mathfrak{u}(N).$ (15)
The induced curve then $\gamma(t)=e^{t\widetilde{\eta}}U$ satisfies $\gamma(0)=U$ and $\dot{\gamma}(0)=\widetilde{\eta}U=\eta$, verifying that $\mathrm{R}_{U}^{\text{exp}}$ is a valid retraction. Other examples are provided in Appendix˜A.
At this point, we can consider the simplest first-order optimization algorithm, i.e., the Riemannian gradient ascent method with the retraction $\mathrm{R}=\mathrm{R}^{\text{exp}}$ defined in Eq.˜15. Its update rule takes the form ($k=0,1,\ldots$)
$U_{k+1}=\mathrm{R}_{U_{k}}^{\text{exp}}\!\left(t_{k}\operatorname{grad}f(U_{k})\right)=e^{\,t_{k}[H,\psi_{k}]}\,U_{k},$ (16)
where $\psi_{k}:=U_{k}\psi_{0}U_{k}^{\dagger}$ denotes the intermediate quantum state after the $k$-th iteration. Setting the initial gate to $U_{0}=I$, the resulting state after $T$ iterations is
$|\psi_{T}\rangle=e^{t_{T-1}[H,\psi_{T-1}]}\cdots e^{t_{1}[H,\psi_{1}]}e^{t_{0}[H,\psi_{0}]}|\psi_{0}\rangle.$ (17)
However, exactly implementing the newly introduced gate $e^{t_{k}\left[H,\psi_{k}\right]}$ on quantum hardware is very challenging. Current works [40, 41, 48, 32] address this issue by using stochastic subspace projected gradient. In these approaches, the Riemannian gradient $\left[H,\psi_{k}\right]$ is projected onto a small number of randomly selected Pauli words, which form a basis of a random subspace in the tangent space $\mathfrak{u}(N)$; a Trotter approximation [51] is then used as the retraction (see Eq.˜55 in Appendix˜A). However, such methods can only exploit partial information of the Riemannian gradient at each iteration. As we will soon see, for the Hamiltonian $H=H^{2}$ arising in the Grover setting, there exists an exact implementation of the Riemannian gradient, rather than merely an approximate one.
III.2 Grover-compatible retractions
The difficulty of implementing the gate $e^{t_{k}[H,\psi_{k}]}$ in Eq.˜16 suggests that the Riemannian exponential map $\mathrm{R}_{U}=\mathrm{R}^{\text{exp}}_{U}$ in Eq.˜15 may not be a practical choice of retraction. Note that the convergence guarantees of manifold optimization [1, 6] only require a retraction $\mathrm{R}_{U}$ to satisfy the first-order properties: $\gamma(0)=U$, $\dot{\gamma}(0)=\eta$ in Eq.˜14; while its specific form is unrestricted. This motivates us to adopt another retraction that is easier to implement on quantum devices.
Throughout the remainder of the paper, let $H=H^{\dagger}=H^{2}$ be an orthogonal projector on Hilbert space $\mathcal{H}$ and $M=\operatorname{rank}(H)$; let $\left|\psi_{0}\right\rangle\in\mathcal{H}$ be a unit vector and set $\psi_{0}=\left|\psi_{0}\right\rangle\left\langle\psi_{0}\right|$. Assume that $\left\langle\psi_{0}\right|H\left|\psi_{0}\right\rangle\notin\{0,1\}$. Define the Frobenius norm by $\|A\|_{F}^{2}=\langle A,A\rangle$. We begin with an auxiliary lemma.
Lemma 1 ([31, Lemma 3.2]).
Let $\psi=|\psi\rangle\langle\psi|$ be any pure state and define $q:=\langle\psi|H|\psi\rangle$. Define two skew-Hermitian operators by $X:=[H,\psi]$ and $Y:=i[H,X]$. Then $\|X\|_{F}=\|Y\|_{F}=\sqrt{2q(1-q)}$ and $\langle X,Y\rangle=0$.
The following Theorem˜1 shows that if the circuit structure consists only of the diffusion and oracle operators defined in Eqs.˜2 and 3, then the output states always remain in the Grover plane. More importantly, the corresponding Riemannian gradient also stays within a fixed two-dimensional subspace of the tangent space. Note that $X_{0}$ and $Y_{0}$ defined in Eq.˜19 are orthogonal, as stated in Lemma˜1.
Theorem 1 (Invariant 2D gradient subspace [31, Theorem 3.3]).
Initialize the unitary by setting $U_{0}=I$. For $k=0,1,\ldots$, consider the update rule:
[…]
and define $U_{k+1}:=V_{k}U_{k}$. Moreover, let $\psi_{k}:=\left|\psi_{k}\right\rangle\left\langle\psi_{k}\right|$, where $\left|\psi_{k}\right\rangle=U_{k}\left|\psi_{0}\right\rangle$; equivalently, $\left|\psi_{k+1}\right\rangle=V_{k}\left|\psi_{k}\right\rangle$. Define the scalars $q_{k}:=f(U_{k})=\left\langle\psi_{k}\right|H\left|\psi_{k}\right\rangle\in[0,1],$ and the skew-Hermitian operators
$X_{0}:=\left[H,\psi_{0}\right],\quad Y_{0}:=i\left[H,X_{0}\right].$ (19)
Then, for all $k\geq 0$, the following statements hold:
  1. 1.
    The state $|\psi_{k}\rangle$ remains in the fixed 2-dimensional subspace (called Grover plane)
    $|\psi_{k}\rangle\in\mathcal{S}:=\operatorname{span}_{\mathbb{C}}\{|\psi_{0}\rangle,H|\psi_{0}\rangle\}\subseteq\mathcal{H}.$ (20)
  2. 2.
    The skew-Hermitian part of Riemannian gradient, $H,\psi_{k}]$, remains in the fixed 2-dimensional real subspace
    $H,\psi_{k}]\in\mathcal{W}:=\operatorname{span}_{\mathbb{R}}\{X_{0},Y_{0}\}\subseteq\mathfrak{u}(N),$ (21)
    and $\left\|\left[H,\psi_{k}\right]\right\|_{F}=\sqrt{2q_{k}\left(1-q_{k}\right)}$.
Theorem˜1 provides an important insight: if the current gradient [^1] In what follows, whenever there is no ambiguity, the term gradient also refers to the skew-Hermitian part of the Riemannian gradient. lies in the 2-dimensional subspace $\mathcal{W}\subseteq\mathfrak{u}(N)$ (when $U_{0}=I$, the initial Riemannian gradient is simply $X_{0}\in\mathcal{W}$), then it suffices to consider a retraction that is well-defined only on $\mathcal{W}$, rather than on the full $N^{2}=4^{n}$-dimensional space $\mathfrak{u}(N)$ [^2] A typical contrast is the Trotter retraction defined in Eq. 55 (see Appendix A).. First, to ensure that it is a valid retraction, it must satisfy the standard first-order properties in Eq.˜14. Second, to guarantee that the next Riemannian gradient remains in $\mathcal{W}$, its construction must involve only the form $e^{i\theta H}$ and $e^{i\theta\psi_{0}}$. In this way, the whole procedure forms a closed loop. This naturally leads to the following definition.
Definition 1.
A mapping $\mathrm{R}_{U}$ is called a Grover-compatible retraction if for all $U\in\mathrm{U}(N)$ the following conditions hold:
  1. 1.
    $\mathrm{R}_{U}:\mathcal{W}U\rightarrow\mathrm{U}(N)$ is a well-defined retraction restricted to the domain $\mathcal{W}U\subseteq T_{U}=\mathfrak{u}(N)U$; that is, the induced curve $\gamma(t)=\mathrm{R}_{U}(t\eta)$ for $t\geq 0$ satisfies
    $\gamma(0)=U,\quad\dot{\gamma}(0)=\eta,\quad\forall\eta\in\mathcal{W}U.$ (22)
  2. 2.
    $\mathrm{R}_{U}$ admits the form $\mathrm{R}_{U}(\eta)=V(\eta)U,$ where $V(\eta)$ is a finite product of the form $e^{i\theta H}$ and $e^{i\theta\psi_{0}}$.
In fact, retractions satisfying the above definition are easy to construct, and their forms may vary. Here we introduce a $5$-factor retraction in Example˜1. According to the number $K$ of factors of the form $e^{i\theta H}$ and $e^{i\theta\psi_{0}}$ it contains, we refer to it as a $K$-factor retraction. More examples, such as $6$-factor and $8$-factor retractions, can be found in [31, Proposition 3.5].
Example 1 ($5$-factor retraction).
For any $U\in\mathrm{U}(N)$ and $\eta\in\mathcal{W}U$, write the skew-Hermitian part $\widetilde{\eta}:=\eta U^{\dagger}\in\mathcal{W}$ in the form $\widetilde{\eta}=xX_{0}+yY_{0}$ for unique coefficients $(x,y)\in\mathbb{R}^{2}$. Let $A:=\operatorname{atan2}(y,x)$ and $R:=\sqrt{x^{2}+y^{2}}$, which correspond to the argument and modulus of the complex number $x+iy$, and define
$a_{1}:=A+\frac{\pi}{2},\;a_{2}:=A-\frac{\pi}{2},\;b_{1}:=-\frac{R}{2},\;b_{2}:=\frac{R}{2}.$ (23)
The $5$-factor retraction is defined by
$\mathrm{R}^{(5)}_{U}(\eta):=e^{ia_{1}H}e^{ib_{1}\psi_{0}}e^{i(a_{2}-a_{1})H}e^{ib_{2}\psi_{0}}e^{-ia_{2}H}\,U.$ (24)
Since $a_{2}-a_{1}=-\pi$, the middle $H$-exponential simplifies to $e^{-i\pi H}$. For $t\geq 0$, the associated curve is given by
$\gamma(t)=\underbrace{e^{ia_{1}H}e^{itb_{1}\psi_{0}}e^{-i\pi H}e^{itb_{2}\psi_{0}}e^{-ia_{2}H}}_{\text{newly added gates }V(t;x,y):=}\,U.$ (25)
It can be verified that $\gamma(0)=U$ and $\dot{\gamma}(0)=\eta\in\mathcal{W}U$.
Indeed, the Grover-compatible retractions $\mathrm{R}_{U}\colon\mathcal{W}U\simeq\mathbb{R}^{2}\rightarrow\mathrm{U}(N)$ admit the following general abstract form:
$\mathrm{R}_{U}(t\eta)=\underbrace{\left(\prod_{\ell=1}^{K}e^{i\,\theta_{\ell}^{(1)}(t;x,y)\,H}\,e^{i\,\theta_{\ell}^{(2)}(t;x,y)\,\psi_{0}}\right)}_{\text{newly added gates }V(t;x,y):=}\,U,$ (26)
for $\widetilde{\eta}:=\eta U^{\dagger}=xX_{0}+yY_{0}\in\mathcal{W}$. The order of $H$- and $\psi_{0}$-exponentials does not matter. Now, we provide the Riemannian gradient ascent (RGA) method employing Grover-compatible retractions in Algorithm˜1.
1
Input: Choose a Grover-compatible retraction $\mathrm{R}$ in Definition˜1 with formula Eq.˜26, initial point $U_{0}=I$, and tolerance $\varepsilon$.
2
3Set $k:=0$, $q_{0}:=M/N$;
4
5while $\|[H,\psi_{k}]\|>\varepsilon$ do
6   Decompose $H,\psi_{k}]=x_{k}X_{0}+y_{k}Y_{0}$ to find $(x_{k},y_{k})\in\mathbb{R}^{2}$;
7
8  Choose a step size $t_{k}>0$;
9
10  Update $U_{k+1}:=V(t_{k};x_{k},y_{k})U_{k}$ by Eq.˜26;
11
12  Update $|\psi_{k+1}\rangle:=V(t_{k};x_{k},y_{k})|\psi_{k}\rangle$;
13
14  Compute new cost value $q_{k+1}:=\langle\psi_{k+1}|H|\psi_{k+1}\rangle$;
15
16  Compute $\|[H,\psi_{k+1}]\|_{F}:=\sqrt{2q_{k+1}(1-q_{k+1})}$;
17
18  Set $k:=k+1$;
19
Algorithm 1 Grover-compatible Riemannian gradient ascent (RGA) method
[31, Theorem 4.6] provides explicit results on the convergence and complexity of Algorithm˜1. Suppose we run Algorithm˜1 with the 5-factor retraction in Eq.˜24, and choose a fixed step size $t_{k}=1/L_{\text{Rie}}$, where $L_{\text{Rie}}:=2+N/\sqrt{2M(N-M)}\in\mathcal{O}(\sqrt{N/M})$. Then, for any $0<\varepsilon\leq q_{0}$, the iterates satisfy $1-q_{T}\leq\varepsilon$ within at most $T=\left\lceil 6L_{\text{Rie}}\log\left(1/\varepsilon\right)\right\rceil$ iterations. In summary, we establish an iteration complexity of $\mathcal{O}(\sqrt{N/M}\log(1/\varepsilon))$ to reach an $\varepsilon$-global maximizer, which is consistent with Grover’s quadratic speedup. The term $\log(1/\varepsilon)$ indicates that the method converges linearly with respect to the accuracy $\varepsilon$.
Remark 2.
Note that as long as the initial state $\left|\psi_{0}\right\rangle$ is the uniform state, Algorithm˜1 is guaranteed to converge to the desired target state $\left|\psi^{\star}\right\rangle$ defined in Eq.˜4. Recall that $\left|\psi^{\star}\right\rangle=\left(1/\sqrt{q_{0}}\right)H\left|\psi_{0}\right\rangle$. For any $\left|\psi_{k}\right\rangle\in\mathcal{S}=\operatorname{span}_{\mathbb{C}}\left\{\left|\psi_{0}\right\rangle,H\left|\psi_{0}\right\rangle\right\}$, the cost value satisfies $q_{k}=\left\langle\psi_{k}\right|H\left|\psi_{k}\right\rangle=\left|\left\langle\psi_{k}|\psi^{\star}\right\rangle\right|^{2}$. Therefore, when $q_{k}\rightarrow 1$, the states $\left|\psi_{k}\right\rangle$ and $\left|\psi^{\star}\right\rangle$ differ only by a global phase.
III.3 Classical simulability of Algorithm˜1
Algorithm˜1 does not need to be executed directly on a quantum device. As shown in Theorem˜2, the procedure is classically simulable. Intuitively, the entire evolution of Algorithm˜1 takes place within a two-dimensional subspace. We choose $u:=H\left|\psi_{0}\right\rangle$ and $v:=(I-H)\left|\psi_{0}\right\rangle$ as a basis for the Grover plane. It therefore suffices to represent the state $\left|\psi_{k}\right\rangle$ and the gradient $\left[H,\psi_{k}\right]$ in this basis. The detailed proof can be found in [31, Theorem 3.8]. In practice, the gate angles can be computed in advance on a classical computer, after which the resulting circuit is executed on quantum hardware.
Theorem 2 (Classical simulability of Algorithm˜1 [31, Theorem 3.8]).
Consider the iterative process in Algorithm˜1. With the initial triplet $(x_{0},y_{0},q_{0}):=(1,0,q_{0})$ and the $2\times 2$ matrix $\Psi_{0}=\begin{bmatrix}q_{0}&1-q_{0}\\ q_{0}&1-q_{0}\end{bmatrix}$, there exists an explicit, classically computable process
$F_{\mathrm{RGA}}:(x_{k},y_{k},q_{k};t_{k})\mapsto(x_{k+1},y_{k+1},q_{k+1}),$ (27)
described as follows. Initialize $\alpha_{0}=\beta_{0}=1$, and $z_{0}=\alpha_{0}\bar{\beta}_{0}.$ For each $k=0,1,\ldots,$
  1. 1.
    For the additional update gates $V(t_{k};x_{k},y_{k})$ in Eq.˜26, define the corresponding $2\times 2$ matrix
    $M_{k}:=\prod_{\ell=1}^{K}E_{H}\!\big(\theta_{\ell}^{(1)}(t_{k};x_{k},y_{k})\big)\,E_{\psi_{0}}\!\big(\theta_{\ell}^{(2)}(t_{k};x_{k},y_{k})\big),$
    where $E_{H}(\theta)=\begin{bmatrix}e^{i\theta}&0\\ 0&1\end{bmatrix}$, $E_{\psi_{0}}(\theta)=I_{2}+(e^{i\theta}-1)\Psi_{0}$.
  2. 2.
    Update $\begin{pmatrix}\alpha_{k+1}\\ \beta_{k+1}\end{pmatrix}=M_{k}\begin{pmatrix}\alpha_{k}\\ \beta_{k}\end{pmatrix}$, and let $q_{k+1}=q_{0}|\alpha_{k+1}|^{2},$ $z_{k+1}=\alpha_{k+1}\bar{\beta}_{k+1},$ $x_{k+1}=\Re(z_{k+1}),$ $y_{k+1}=\Im(z_{k+1})$.
  3. 3.
    Set $k\leftarrow k+1$ and return to step 1.
As we will see later, the Riemannian Newton method proposed in this work is also classically simulable and can be obtained with modifications based on Theorem˜2.
IV Grover-compatible Riemannian modified Newton method
In the classical Euclidean setting, consider the optimization problem $\max_{x\in\mathbb{R}^{n}}f(x)$. A prototypical second-order algorithm is the Newton method [10, 28, 46], which updates via the Hessian matrix:
$x_{k+1}=x_{k}+d_{k},\quad d_{k}=-\left[\nabla^{2}f\left(x_{k}\right)\right]^{-1}\nabla f\left(x_{k}\right),$ (28)
where $d_{k}$ denotes the Newton direction. Under standard regularity conditions, the method exhibits quadratic convergence in a neighborhood of the solution. The Newton method has also been extended to optimization on manifolds [2, 1, 14], while preserving the same quadratic convergence rate with respect to the accuracy $\varepsilon$.
In this section, building on the first-order framework developed in the previous section, we will replace the Riemannian gradient direction with the Riemannian Newton direction, thereby obtaining a Grover-compatible Riemannian Newton method. This modification is expected to improve the complexity from $\mathcal{O}(\sqrt{N/M}\log(1/\varepsilon))$ to $\mathcal{O}(\sqrt{N/M}\log\log(1/\varepsilon))$.
IV.1 Riemannian Hessian
For $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$, the Hessian matrix $\nabla^{2}f(x)$ at a point $x$ captures the second-order information of $f$. It is an $n\times n$ symmetric matrix and can be viewed as a self-adjoint linear operator on $\mathbb{R}^{n}$. By analogy, when $f$ is defined on a manifold, the Riemannian Hessian (operator) can be regarded as a self-adjoint linear operator acting on the tangent space. For the cost function defined in Eq.˜7, we have the following lemma.
Lemma 2 ([32, Proposition 1]).
For the cost function $f:\mathrm{U}(N)\rightarrow\mathbb{R}$ defined by $f(U)=\operatorname{Tr}(HU\psi_{0}U^{\dagger})$, the Riemannian Hessian of $f$ at $U\in\mathrm{U}(N)$ is the self-adjoint linear map $\operatorname{Hess}f(U):T_{U}\rightarrow T_{U}$,
$\operatorname{Hess}f(U)[\Omega U]=\frac{1}{2}\big([H,[\Omega,\psi_{U}]]+[[H,\Omega],\psi_{U}]\big)U,$ (29)
where $\psi_{U}=U\psi_{0}U^{\dagger}$ and $T_{U}=\mathfrak{u}(N)U$. Identifying $T_{U}\simeq\mathfrak{u}(N)$ yields $\widetilde{\operatorname{Hess}}f(U):\mathfrak{u}(N)\rightarrow\mathfrak{u}(N),$
$\widetilde{\operatorname{Hess}}f(U)[\Omega]=\frac{1}{2}\big([H,[\Omega,\psi_{U}]]+[[H,\Omega],\psi_{U}]\big),$ (30)
which is self-adjoint on $\mathfrak{u}(N)$.
The self-adjointness means that $\langle\Xi,\widetilde{\operatorname{Hess}}f(U)[\Omega]\rangle=\langle\widetilde{\operatorname{Hess}}f(U)[\Xi],\Omega\rangle$ for any $\Omega,\Xi\in\mathfrak{u}(N)$. A derivation of this Hessian expression can be found in [32, Proposition 1]. An equivalent definition is based on the second-order Taylor expansion on the manifold:
$f\left(\mathrm{R}_{U}(t\eta)\right)$ (31)
$=f(U)+t\langle\operatorname{grad}f(U),\eta\rangle+\frac{t^{2}}{2}\langle\eta,\operatorname{Hess}f(U)[\eta]\rangle+\mathcal{O}(t^{3}),$ (32)
where $\mathrm{R}_{U}$ is a retraction and $\eta\in T_{U}$. In Appendix˜A, we verify that Eq.˜29 ensures the above expansion holds.
In the classical Euclidean setting, the Newton direction $d_{k}$ is defined as the solution to the Newton equation:
$\nabla^{2}f\left(x_{k}\right)d_{k}=-\nabla f\left(x_{k}\right).$ (33)
By analogy, the Riemannian Newton direction $\Omega_{k}^{\mathrm{N}}U_{k}\in T_{U_{k}}$ is defined as the solution to the Riemannian Newton equation: $\operatorname{Hess}f(U_{k})[\Omega_{k}^{\text{N}}U_{k}]=-\operatorname{grad}f(U_{k})$, i.e.,
$\widetilde{\operatorname{Hess}}f(U_{k})[\Omega_{k}^{\text{N}}]=-\widetilde{\operatorname{grad}}f(U_{k}).$ (34)
Solving these equations is generally challenging. In [32], Pauli words are used as a basis of $\mathfrak{u}(N)$ to convert the above operator equation into a matrix equation. However, this approach is computationally expensive.
In contrast, in the Grover setting, where the Hamiltonian satisfies $H=H^{2}$, a remarkable relationship arises: the Riemannian gradient is always an eigenvector of the Riemannian Hessian. Consequently, the Newton equation admits an explicit solution. This result is formalized in the following theorem, with the proof provided in Appendix˜B.
Theorem 3.
Let $H=H^{\dagger}=H^{2}$ be an orthogonal projector, and let $\psi=|\psi\rangle\langle\psi|$ be an arbitrary pure state. Define $q:=\operatorname{Tr}(H\psi)\in[0,1]$ and $g:=[H,\psi]$. Consider the map $L(\Omega):=\tfrac{1}{2}\bigl([H,[\Omega,\psi]]+[[H,\Omega],\psi]\bigr)$. Then L(g)=(12q)gL(g)=(1-2q)g.
Note that the above result holds only under the condition $H=H^{2}$, and does not hold in general.
IV.2 Grover-compatible Riemannian modified Newton method
In this subsection, we are now ready to present a Riemannian Newton method for problem Eq.˜5, which is both Grover-compatible and classically simulable, similar to Algorithm˜1. Let $q_{k}=f\left(U_{k}\right)$, and define
$\mathrm{g}_{k}:=\widetilde{\operatorname{grad}}f\left(U_{k}\right)\neq 0,\quad\mathrm{H}_{k}:=\widetilde{\operatorname{Hess}}f\left(U_{k}\right).$ (35)
Then Theorem˜3 shows that
$\mathrm{H}_{k}\left[\mathrm{g}_{k}\right]=\lambda_{k}\mathrm{g}_{k},\quad\lambda_{k}:=1-2q_{k}.$ (36)
Since $q_{k}\in[0,1]$, the eigenvalue $\lambda_{k}$ lies in the interval [1,1][-1,1]. The original Newton equation $\mathrm{H}_{k}\left[\Omega_{k}^{\mathrm{N}}\right]=-\mathrm{g}_{k}$ in Eq.˜34 therefore admits the explicit solution
$\Omega_{k}^{\mathrm{N}}=\frac{1}{-\lambda_{k}}\mathrm{g}_{k}\in\mathcal{W},$ (37)
whenever $q_{k}\neq\frac{1}{2}$. (Recall that $\mathcal{W}:=\operatorname{span}_{\mathbb{R}}\left\{X_{0},Y_{0}\right\}\subseteq\mathfrak{u}(N)$ as defined in Theorem˜1.) The corresponding standard Newton update is then given by $U_{k+1}=\mathrm{R}_{U_{k}}(\Omega_{k}^{\mathrm{N}})$ with unit step size $t_{k}=1.$
Remark 3.
Since the Riemannian gradient is always an eigenvector of the Riemannian Hessian, the Newton direction is collinear with the gradient. Then, the Newton method can be viewed as a scaled gradient method.
However, the standard Newton method converges only when the initial point $U_{0}$ is sufficiently close to the solution. To improve numerical stability, we therefore employ a modified Newton method [46]. Specifically, we introduce a parameter $\mu_{k}\geq 0$ to modify the Hessian operator and solve the modified Newton equation
$\left(\mathrm{H}_{k}-\mu_{k}\mathrm{I}\right)\left[\Omega_{k}^{\mathrm{N}}\right]=-\mathrm{g}_{k},$ (38)
where $\mathrm{I}$ denotes the identity operator on the tangent space $\mathfrak{u}(N)$. By Theorem˜3, this equation also admits a solution, namely the modified Newton direction
$\Omega_{k}^{\mathrm{N}}=\gamma_{k}\mathrm{g}_{k}\in\mathcal{W},\quad\gamma_{k}:=\frac{1}{\mu_{k}-\lambda_{k}},$ (39)
whenever $\mu_{k}\neq\lambda_{k}$. If we set $\mu_{k}>\lambda_{k}$, then
$\left\langle\Omega_{k}^{\mathrm{N}},\mathrm{g}_{k}\right\rangle=\gamma_{k}\left\|\mathrm{g}_{k}\right\|^{2}=\gamma_{k}\,2q_{k}\left(1-q_{k}\right)>0.$ (40)
Hence, the modified Newton direction is an ascent direction, and there always exists a step size $t_{k}>0$ along this direction that increases the cost value. The parameter $\mu_{k}$ is determined according to the following two cases:
  1. 1.
    If $q_{k}>0.5$, then $\lambda_{k}<0$. In this case, the Hessian operator is already negative along the gradient direction. We therefore set $\mu_{k}=0$, which yields $\gamma_{k}=\frac{1}{-\lambda_{k}}$ like Eq.˜37.
  2. 2.
    If $q_{k}\leq 0.5$, then $\lambda_{k}\geq 0$. In this case, the Hessian operator is positive along the gradient direction, and the unmodified Newton step would point toward a descent direction. To avoid this, we set $\mu_{k}=\lambda_{k}+\delta$ with $\delta>0$, which yields $\gamma_{k}=\frac{1}{\delta}$.
Combining the two cases above, we set the scaling factor
$\gamma_{k}=\frac{1}{\max\left(\delta,2q_{k}-1\right)}$ (41)
where $\delta>0$ is a small constant, typically chosen in the range $10^{-3}\sim 10^{-6}$.
After ensuring that the Newton direction $\Omega_{k}^{\mathrm{N}}$ is an ascent direction, we further employ an Armijo backtracking line search; see [46, Algorithm 3.2]. Specifically, we seek the smallest integer $m\geq 0$ such that $t_{k}=\rho^{m}$ satisfies
$f\left(\mathrm{R}_{U_{k}}(t_{k}\Omega_{k}^{\mathrm{N}})\right)\geq f\left(U_{k}\right)+c\,t_{k}\left\langle\Omega_{k}^{\mathrm{N}},\mathrm{g}_{k}\right\rangle,$ (42)
where $\rho=0.5$ and $c=10^{-4}$. Note that $\left\langle\Omega_{k}^{\mathrm{N}},\mathrm{g}_{k}\right\rangle>0$, which guarantees that the above procedure terminates after finitely many iterations; see [46, Lemma 3.1]. Consequently, the cost value increases monotonically. Finally, a complete description of the proposed algorithm is given in Algorithm˜2.
This modified Newton method with line search ensures convergence from any initial point (i.e., it enjoys the global convergence property of first-order gradient methods), while achieving quadratic convergence once the iterate is sufficiently close to the solution. Intuitively, during the early iterations, Algorithm˜2 behaves similarly to Algorithm˜1. When the iterate is sufficiently close to the solution (so that $q_{k}>0.5$ and no direction modification is required), the Armijo condition is always satisfied with $t_{k}=1$; see [46, Theorem 3.6 & 3.8]. Consequently, Algorithm˜2 reduces to the standard Newton method.
1
Input: Choose a Grover-compatible retraction $\mathrm{R}$ in Definition˜1 with formula Eq.˜26, initial point $U_{0}=I$, damping $\delta=10^{-3}$, Armijo parameters $c=10^{-4}$, $\rho=\tfrac{1}{2}$, and tolerance $\varepsilon$.
2
3Set $k:=0$, $q_{0}:=M/N$, and initial squared gradient norm $G_{0}:=2q_{0}(1-q_{0})$;
4
5while $\sqrt{G_{k}}>\varepsilon$ do
6
7  Decompose $H,\psi_{k}]=x_{k}X_{0}+y_{k}Y_{0}$ to find $(x_{k},y_{k})\in\mathbb{R}^{2}$;
8
9  Set modified Newton scaling $\gamma_{k}:=\frac{1}{\max(\delta,\,2q_{k}-1)}$;
10
11  Set step size $t:=1$;
12 repeat
13      Set $|\psi_{\mathrm{trial}}\rangle:=V(t\gamma_{k};x_{k},y_{k})|\psi_{k}\rangle$ by Eq.˜26;
14      Compute $q_{\mathrm{trial}}:=\langle\psi_{\mathrm{trial}}|H|\psi_{\mathrm{trial}}\rangle$;
15    if $q_{\mathrm{trial}}<q_{k}+c\,t\,\gamma_{k}\,G_{k}$ then
16         Set $t:=\rho t$;
17
18
19 until $q_{\mathrm{trial}}\geq q_{k}+c\,t\,\gamma_{k}\,G_{k}$;
20  Update $U_{k+1}:=V(t\gamma_{k};x_{k},y_{k})U_{k}$;
21   Update $|\psi_{k+1}\rangle:=|\psi_{\mathrm{trial}}\rangle$;
22   Update $q_{k+1}:=q_{\mathrm{trial}}$;
23   Update $G_{k+1}:=2q_{k+1}(1-q_{k+1})$;
24   Set $k:=k+1$;
25
Algorithm 2 Grover-compatible Riemannian modified Newton (RMN) method
IV.3 Classical simulability of Algorithm˜2
Compared with the gradient ascent method in Algorithm˜1, the modified Newton method in Algorithm˜2 differs only by the introduction of a scaling factor $\gamma_{k}$ and a line-search procedure. Building on the classical simulability of Algorithm˜1 established in Theorem˜2, we can directly derive a classically simulable version of Algorithm˜2, as stated in the following theorem. All operations reduce to computations with $2\times 2$ matrices.
Theorem 4 (Classical simulability of Algorithm˜2).
Consider the iterative process in Algorithm˜2. With the initial triplet $(x_{0},y_{0},q_{0}):=(1,0,q_{0})$ and the $2\times 2$ matrix $\Psi_{0}=\begin{bmatrix}q_{0}&1-q_{0}\\ q_{0}&1-q_{0}\end{bmatrix}$, there exists an explicit, classically computable process
$F_{\mathrm{RMN}}:(x_{k},y_{k},q_{k})\mapsto(x_{k+1},y_{k+1},q_{k+1}),$ (43)
described as follows. Initialize $G_{0}=2q_{0}(1-q_{0}),$ $\alpha_{0}=\beta_{0}=1$, and $z_{0}=\alpha_{0}\bar{\beta}_{0}.$ For each $k=0,1,\dots$,
  1. 1.
    Set modified scaling $\gamma_{k}:=\tfrac{1}{\max(\delta,\,2q_{k}-1)}.$
  2. 2.
    For the trial step $t$, and additional update gates $V(t\gamma_{k};x_{k},y_{k})$ in Eq.˜26, define the corresponding $2\times 2$ matrix
    $M_{k}(t):=\prod_{\ell=1}^{K}E_{H}\!\big(\theta_{\ell}^{(1)}(t\gamma_{k};x_{k},y_{k})\big)\,E_{\psi_{0}}\!\big(\theta_{\ell}^{(2)}(t\gamma_{k};x_{k},y_{k})\big),$
    where $E_{H}(\theta)=\begin{bmatrix}e^{i\theta}&0\\ 0&1\end{bmatrix}$, $E_{\psi_{0}}(\theta)=I_{2}+(e^{i\theta}-1)\Psi_{0}$.
  3. 3.
    For the trial step $t$, define $\begin{pmatrix}\alpha_{k}^{\mathrm{temp}}(t)\\[5.69054pt] \beta_{k}^{\mathrm{temp}}(t)\end{pmatrix}:=M_{k}(t)\begin{pmatrix}\alpha_{k}\\[5.69054pt] \beta_{k}\end{pmatrix}$, and $q_{k}^{\mathrm{temp}}(t):=q_{0}\big|\alpha_{k}^{\mathrm{temp}}(t)\big|^{2}.$
  4. 4.
    Let $m_{k}\geq 0$ be the smallest nonnegative integer such that the Armijo condition
    $q_{k}^{\mathrm{temp}}(\rho^{m_{k}})\;\geq\;q_{k}+c\,\rho^{m_{k}}\gamma_{k}\,G_{k}$ (44)
    holds, and set $t_{k}:=\rho^{m_{k}}.$
  5. 5.
    Update $\begin{pmatrix}\alpha_{k+1}\\[5.69054pt] \beta_{k+1}\end{pmatrix}=M_{k}(t_{k})\begin{pmatrix}\alpha_{k}\\[5.69054pt] \beta_{k}\end{pmatrix}$, and let $q_{k+1}=q_{0}|\alpha_{k+1}|^{2}$, $G_{k+1}=2q_{k+1}(1-q_{k+1})$, $z_{k+1}=\alpha_{k+1}\bar{\beta}_{k+1}$, $x_{k+1}=\Re(z_{k+1})$, $y_{k+1}=\Im(z_{k+1})$.
  6. 6.
    Set $k\leftarrow k+1$ and return to step 1.
V Numerical experiments
In this section, we compare the following methods via numerical simulations:
  • the Grover-compatible Riemannian gradient ascent (RGA) method (Algorithm˜1) [31], and
  • the proposed Grover-compatible Riemannian modified Newton (RMN) method (Algorithm˜2).
All experiments are implemented in the NumPy package and conducted on a MacBook Pro (2021) with an Apple M1 Pro processor and 16 GB of memory.
Consider an unstructured search problem of size $N=2^{n}$ for an $n$-qubit system, with a single marked item ($M=1$). We always use 5-factor retraction defined in Example˜1. The initial state $|\psi_{0}\rangle$ is the uniform superposition. The progress of the algorithms is measured by the cost value $q_{k}=f\left(U_{k}\right)=\operatorname{Tr}\left(H\psi_{k}\right),$ which represents the success probability of identifying the marked item and should converge to 1 under RGA and RMN. Given $\varepsilon>0$ (typically $10^{-10}$), the algorithms terminate when $\left|1-q_{k}\right|<\varepsilon$.
Correctness of classical simulability.
We first examine whether the classical procedures described in Theorems˜2 and 4 can faithfully simulate the quantum implementations of Algorithms˜1 and 2. To this end, we set $n=4$ qubits. At each iteration, we evaluate three quantities: the cost value $q_{k}$, and the coefficients $x_{k},y_{k}$. Recall that $x_{k}$ and $y_{k}$ are the expansion coefficients of the Riemannian gradient $\left[H,\psi_{k}\right]$ on the two-dimensional subspace $\mathcal{W}$ with respect to the basis $\left\{X_{0},Y_{0}\right\}$.
Fig.˜2 reports the absolute errors of these three quantities for RGA ((a)–(c)) and RMN ((d)–(f)), with and without classical simulation. Here, “without classical simulation” refers to a direct NumPy implementation using explicit $N\times N$ full matrix operations, following Algorithms˜1 and 2 in a brute-force manner. The results show that the classical simulation accurately reproduces the true iterations up to machine precision $(10^{-16})$. Therefore, all subsequent experiments are conducted under classical simulation.
Figure 2: Absolute errors of the cost value $q_{k}$ and expansion coefficients $x_{k},y_{k}$ between the classical simulation and the explicit full matrix implementation. (a)–(c) display the results for RGA, and (d)–(f) for RMN. All errors remain around machine precision ($10^{-16}$), verifying that the classical procedures accurately simulate the algorithms.
Comparison of convergence rates.
As a first-order optimization method, RGA is shown to achieve asymptotic linear convergence with respect to the error [31]. Let the function value error be denoted as $\delta_{k}:=\left|1-q_{k}\right|$. Then, there exist an integer $k_{1}$ and a contraction factor $\rho\in(0,1)$ such that $\delta_{k+1}\leq\rho\delta_{k},\forall k\geq k_{1}.$ In contrast, the proposed RMN is a second-order method. RMN achieves local quadratic convergence: there exist an integer $k_{2}$ and a constant $C>0$ such that $\delta_{k+1}\leq C\delta_{k}^{2},\forall k\geq k_{2}.$
It is well known that quadratic convergence is significantly faster than linear convergence. For each case, we run both RGA (with fixed step size 0.5) and RMN. Fig.˜3 shows the results on a log scale, including the gradient norm ((a)–(c)) and the function value error ((d)–(f)). The results show that RGA decreases linearly with respect to the iteration count. In contrast, RMN exhibits quadratic convergence and reaches a high-accuracy solution within only a few iterations. This behavior is evident in both the gradient norm and the function value error. For example, in (d), RMN reduces the error from $10^{-2}$ to $10^{-4}$, and then to $10^{-8}$ in just two iterations.
Figure 3: Convergence comparison between the RGA and RMN methods for problem sizes of $n=5$, 1010, and 1515 qubits. (a)–(c) illustrate the gradient norm, and (d)–(f) show the function value error. The results demonstrate the linear convergence of RGA and the significantly faster, quadratic convergence of RMN.
Scaling with problem size.
RGA is shown in [31] to have complexity $\mathcal{O}(\sqrt{N/M}\log(1/\varepsilon))$. Fixing the accuracy $\varepsilon$ and setting $M=1$, this reduces to the well-known quadratic speedup $\mathcal{O}(\sqrt{N})$, i.e., the iteration count grows linearly with $\sqrt{N}$. The proposed RMN is a second-order optimization method, whose improvement targets the dependence on the accuracy $\varepsilon$ rather than on the problem size NN [^3] Optimization literature typically fixes the problem size and studies how to achieve higher accuracy with fewer iterations.. It is therefore expected to retain the same scaling $\mathcal{O}(\sqrt{N})$ with respect to the problem size. However, due to its quadratic convergence, the overall complexity improves to $\mathcal{O}(\sqrt{N}\log\log(1/\varepsilon)).$
Next, we investigate the dependence of RMN on the problem size $N$. We fix $\varepsilon=10^{-6}$ and vary the number of qubits from $n=2$ to $n=28$, corresponding to $\sqrt{N}$ ranging from 2 to $\sqrt{2^{28}}=16384$. Fig.˜4 shows a clear linear relationship: the number of iterations of RMN to reach the prescribed accuracy scales proportionally to $\sqrt{N}$.
Figure 4: Iteration complexity of RMN versus the square root of the problem size. For a fixed tolerance $\varepsilon$ and varying qubit counts $n\in[2,28]$ ($N=2^{n}$), the required iterations for RMN scale linearly with $\sqrt{N}$. This confirms that the second-order RMN method successfully retains the $\mathcal{O}(\sqrt{N})$ quantum speedup.
VI Discussion
In this work, we develop a deeper understanding of the quantum search problem through the lens of the manifold optimization framework by introducing the Grover-compatible Riemannian modified Newton (RMN) method. While the previous first-order Riemannian gradient ascent (RGA) method [31] successfully recasts quantum search as a maximization problem on the unitary group and achieves $\mathcal{O}(\sqrt{N/M}\log(1/\varepsilon))$ query complexity, it is limited to a linear convergence rate with respect to the target accuracy $\varepsilon$. By incorporating second-order geometric information, we overcome this limitation.
The cornerstone of our RMN method is the theoretical observation that, for projector Hamiltonians $H=H^{2}$, the Riemannian gradient is an eigenvector of the corresponding Riemannian Hessian. This property eliminates the need for costly Hessian inversion, reducing the Newton update to a scaled gradient step. Consequently, our RMN method can achieve a quadratic convergence rate, driving the total complexity down to $\mathcal{O}(\sqrt{N/M}\log\log(1/\varepsilon))$ without incurring additional overhead per iteration. Furthermore, we demonstrate that the classical simulability of the parameter update process is also preserved, allowing for the efficient precomputation of gate angles prior to execution on quantum hardware.
Our results suggest several promising directions for future research. The exact collinearity of the Riemannian gradient and the Newton direction established in our framework (i.e., Theorem˜3) relies strictly on the projective nature of the Grover Hamiltonian ($H=H^{2}$). Therefore, extending this second-order geometric property to more general quantum tasks presents a compelling theoretical challenge. On the other hand, for broader quantum applications such as ground-state preparation for generic molecular Hamiltonians [47, 26], implementing exact Newton or gradient steps may be intractable on quantum computers. In such settings, it is of interest to investigate whether similar invariant gradient subspaces (as in Theorem˜1) can be identified to reduce the effective problem dimensionality. Furthermore, the application of Riemannian quasi-Newton methods [25], particularly Riemannian BFGS [24], could offer a viable pathway to achieving superlinear convergence without the need for explicit Hessian evaluations, enabling accelerations across a wider spectrum of Hamiltonians.
Beyond those theoretical generalizations, future work can evaluate the resilience of the RMN method under realistic hardware noise, as high-precision algorithms may be sensitive to gate infidelities and decoherence in NISQ devices. Finally, integrating this manifold optimization framework with quantum error mitigation techniques may be essential to stabilize its quadratic convergence for practical implementations.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China under the grant numbers 12501419, 12331010 and 12288101, the National Key R&D Program of China under the grant number 2024YFA1012901, the Quantum Science and Technology-National Science and Technology Major Project via Project 2024ZD0301900, and the Fundamental Research Funds for the Central Universities, Peking University. We thank Zongqi Wan for their valuable feedback on the manuscript.
Data availability
The source code of the numerical experiments is publicly available [33].
Appendix A Geometric tools on $\mathrm{U}(N)$
The set of $N\times N$ unitary matrices,
$\mathrm{U}(N)=\left\{U\in\mathbb{C}^{N\times N}\mid U^{\dagger}U=I\right\},$ (45)
forms a Riemannian manifold, despite being a subset of a Euclidean space. As a result, optimization over $\mathrm{U}(N)$ requires geometric tools tailored to the manifold structure. This appendix briefly reviews the basic ingredients for optimization on $\mathrm{U}(N)$, including tangent spaces, Riemannian gradients, retractions, and Riemannian Hessians. For a comprehensive treatment, see [1, 6].
A.1 Tangent space
Geometrically, the tangent space provides a local linear approximation of the manifold at a given point, analogous to a tangent plane to a two-dimensional sphere (see Fig.˜5). From an optimization perspective, it consists of all feasible directions (i.e., tangent vectors) along curves on the manifold. This structure can also be characterized algebraically.
Fix a point $U\in\mathrm{U}(N)$, and consider a smooth curve $t\mapsto U(t)\in\mathrm{U}(N)$ with $U(0)=U$. Differentiating the unitarity constraint $U(t)U(t)^{\dagger}=I$ with respect to $t$ at $t=0$ yields
$U\dot{U}(0)^{\dagger}+\dot{U}(0)U^{\dagger}=0.$ (46)
Letting the tangent vector be $\eta:=\dot{U}(0)$, the above condition implies $(\eta U^{\dagger})^{\dagger}=-(\eta U^{\dagger}).$ Hence, the matrix $\Omega:=\eta U^{\dagger}$ is skew-Hermitian. Consequently, the tangent space at $U$ is defined as
$T_{U}=\left\{\Omega U:\Omega^{\dagger}=-\Omega\right\}=\mathfrak{u}(N)U.$ (47)
Here, $\mathfrak{u}(N)$ denotes the Lie algebra of the unitary group, consisting of all $N\times N$ skew-Hermitian matrices and closed under the Lie bracket [X,Y]=XYYX[X,Y]=XY-YX. In particular, at the identity $I\in\mathrm{U}(N)$, we have $T_{I}=\mathfrak{u}(N)$. Since $\operatorname{dim}\mathfrak{u}(N)=N^{2}$, the manifold $\mathrm{U}(N)$ has dimension $N^{2}$. Note that the tangent spaces of the unitary group differ only through right multiplication by $U$.
Figure 5: Riemannian gradient on the manifold. Viewing $\mathrm{U}(N)$ as a sphere for illustration, the Riemannian gradient $\operatorname{grad}f(U)$ is the orthogonal projection of the Euclidean gradient $\nabla f(U)$ onto the tangent space at $U$. It is perpendicular to the contour and points in the direction of the fastest increase of $f$.
A.2 Riemannian gradient
Consider a function $f:\mathrm{U}(N)\subseteq\mathbb{C}^{N\times N}\rightarrow\mathbb{R}$. The Euclidean gradient $\nabla f(U)$ gives the direction of steepest ascent in the ambient space $\mathbb{C}^{N\times N}$, but it generally does not lie in the tangent space $T_{U}$ and thus is not a feasible direction on the manifold. The standard approach is to orthogonally project the Euclidean gradient onto $T_{U}$ (see [6, Proposition 3.61]); the resulting tangent vector is called the Riemannian gradient, $\operatorname{grad}f(U)\in T_{U}$. Geometrically, the Riemannian gradient is perpendicular to the contour line of $f$ passing through $U$ (see Fig.˜5).
Here, we equip all tangent spaces $T_{U}$ with the Frobenius inner product $\langle A,B\rangle=\operatorname{Tr}\left(A^{\dagger}B\right),$ which is the metric inherited from the ambient space $\mathbb{C}^{N\times N}$. The normal space at $U\in\mathrm{U}(N)$ is defined as the orthogonal complement of the tangent space:
$N_{U}:=\left\{Z\in\mathbb{C}^{N\times N}:\langle Z,X\rangle=0,\forall X\in T_{U}\right\}.$ (48)
Algebraically, this space can be characterized as
$N_{U}=\left\{MU:M^{\dagger}=M\right\}=\mathcal{H}(N)U,$ (49)
where $\mathcal{H}(N)$ is the space of $N\times N$ Hermitian matrices.
Using the orthogonal decomposition $\mathbb{C}^{N\times N}=T_{U}\oplus N_{U}$, any $Z\in\mathbb{C}^{N\times N}$ can be written as $Z=(\Omega+M)U,$ where $\Omega^{\dagger}=-\Omega$ and $M^{\dagger}=M$. Multiplying by $U^{\dagger}$ on the right gives $ZU^{\dagger}=\Omega+M.$ Taking the skew-Hermitian part yields $\Omega=\operatorname{Skew}\left(ZU^{\dagger}\right)$ with $\operatorname{Skew}(A):=\frac{1}{2}(A-A^{\dagger})$. Hence, the orthogonal projection of $Z$ onto $T_{U}$ is
$\mathcal{P}_{U}(Z):=\operatorname{Skew}(ZU^{\dagger})U.$ (50)
The Riemannian gradient is then $\operatorname{grad}f(U)=\mathcal{P}_{U}(\nabla f(U))$.
For the cost function $f(U)=\operatorname{Tr}\left(HU\psi_{0}U^{\dagger}\right)$, the Euclidean gradient is $\nabla f(U)=2HU\psi_{0}$. In general, if $A$ and $B$ are (skew-)Hermitian, then Skew $(2AB)=[A,B]$. Applying this identity with $A:=H$ and $B:=U\psi_{0}U^{\dagger}$, and using Eq.˜50, we obtain
$\operatorname{grad}f(U)=\operatorname{Skew}(2HU\psi_{0}U^{\dagger})U=[H,U\psi_{0}U^{\dagger}]U,$ (51)
We denote the left skew-Hermitian factor by $\widetilde{\operatorname{grad}}f(U):=[H,U\psi_{0}U^{\dagger}]\in\mathfrak{u}(N)$.
A.3 Retractions
Standard additive updates $U+t\eta$ leave the unitary manifold $\mathrm{U}(N)$, as they violate the unitarity constraint. To map such updates back onto $\mathrm{U}(N)$ while preserving the search direction, one introduces a geometric mapping called a retraction [2, 1], which is a central tool in the modern manifold optimization framework.
Define $T\mathrm{U}(N):=\left\{(U,\eta):U\in\mathrm{U}(N),\eta\in T_{U}\right\}.$ A retraction on $\mathrm{U}(N)$ is a smooth mapping
$\mathrm{R}:T\mathrm{U}(N)\rightarrow\mathrm{U}(N),\quad(U,\eta)\mapsto\mathrm{R}_{U}(\eta),$ (52)
such that for any $U\in\mathrm{U}(N)$ and $\eta\in T_{U}$, the induced curve $t\mapsto\gamma(t):=\mathrm{R}_{U}(t\eta)$ satisfies
$\gamma(0)=U\text{ and }\dot{\gamma}(0)=\eta.$ (53)
For the unitary manifold $\mathrm{U}(N)$, several retractions are available. Given $U\in\mathrm{U}(N)$ and a tangent vector $\eta=\tilde{\eta}U\in T_{U}$ with $\tilde{\eta}\in\mathfrak{u}(N)$, common retractions include:
  1. 1.
    Riemannian exponential map [20]:
    $\mathrm{R}_{U}^{\text{exp}}(\eta)=\exp(\widetilde{\eta})\,U,$ (54)
    which generates the exact geodesic curves on $\mathrm{U}(N)$.
  2. 2.
    (First-order) Trotter retraction [51, 36, 32]:
    $\mathrm{R}_{U}^{\text{trt}}(\eta)=\left(\prod_{j=1}^{N^{2}}\exp\!\left(i\omega^{j}P^{j}\right)\right)U,$ (55)
    where $\widetilde{\eta}\in\mathfrak{u}(N)$ is expanded in the $\log_{2}(N)$-qubit Pauli basis $\{P^{j}\}_{j=1}^{N^{2}}$ as $\widetilde{\eta}=i\sum_{j}\omega^{j}P^{j}$ for $\omega^{j}\in\mathbb{R}$.
  3. 3.
    Cayley transform retraction [52]:
    $\mathrm{R}_{U}^{\text{cay}}(\eta)=\left(I-\tfrac{1}{2}\widetilde{\eta}\right)^{-1}\left(I+\tfrac{1}{2}\widetilde{\eta}\right)U.$ (56)
  4. 4.
    QR decomposition retraction [1]:
    $\mathrm{R}_{U}^{\text{qr}}(\eta)=\operatorname{qf}(U+\eta),$ (57)
    where $\operatorname{qf}(A)$ denotes the orthogonal $Q$-factor in the QR decomposition of $A$.
  5. 5.
    Polar decomposition retraction [1]:
    $\mathrm{R}_{U}^{\text{polar}}(\eta)=(U+\eta)(I+\eta^{\dagger}\eta)^{-1/2},$ (58)
    where $(I+\eta^{\dagger}\eta)^{-1/2}$ is the inverse of the unique positive definite Hermitian square root of $I+\eta^{\dagger}\eta$.
The effectiveness of these retractions strongly depends on the computational paradigm. The Riemannian exponential map and the Trotter retraction correspond to physical Hamiltonian evolution, making them naturally suited for implementation on quantum hardware; however, their reliance on matrix exponentials renders them computationally inefficient on classical computers. Conversely, the Cayley, QR, and polar retractions rely on matrix inversions and factorizations. While implementing these nonlinear operations on quantum circuits incurs additional overhead, they are highly optimized within numerical linear algebra, making them the preferred choice in classical computing environments.
A.4 Riemannian Hessian
Unlike the usual setting, where the Hessian is represented by a symmetric matrix of second-order partial derivatives, the Riemannian Hessian at $U$, denoted by $\operatorname{Hess}f(U)$, is a self-adjoint linear operator on $T_{U}$. It describes the covariant derivative of the Riemannian gradient field along a tangent direction, thereby naturally incorporating both the second-order variation of the objective function and the curvature of the manifold.
For the objective $f(U)=\operatorname{Tr}(HU\psi_{0}U^{\dagger})$, evaluating the Riemannian Hessian along a tangent direction $\Omega U\in T_{U}$ yields the explicit formula (see Lemma˜2):
$\operatorname{Hess}f(U)[\Omega U]=\frac{1}{2}\Big([H,[\Omega,\psi]]+[[H,\Omega],\psi]\Big)U,$ (59)
where $\psi:=U\psi_{0}U^{\dagger}$ is the density operator of the current quantum state. This linear operator can be equivalently represented by its action directly on $\mathfrak{u}(N)$:
$\widetilde{\operatorname{Hess}}f(U)[\Omega]=\frac{1}{2}\Big([H,[\Omega,\psi]]+[[H,\Omega],\psi]\Big).$ (60)
A derivation of this Hessian expression can be found in [32, Proposition 1]. To verify the correctness of Eq.˜60, we will apply the second-order Taylor expansion on $\mathrm{U}(N)$. For $\eta=\Omega U\in T_{U}$, using the Riemannian exponential retraction $\mathrm{R}_{U}(\eta)=e^{\Omega}U$ in Eq.˜54, we obtain
$f\left(\mathrm{R}_{U}(t\eta)\right)$
$=f(U)+t\langle\operatorname{grad}f(U),\eta\rangle+\frac{t^{2}}{2}\langle\operatorname{Hess}f(U)[\eta],\eta\rangle+\mathcal{O}(t^{3}),$
$=f(U)+t\langle\widetilde{\operatorname{grad}}f(U),\Omega\rangle+\frac{t^{2}}{2}\langle\widetilde{\operatorname{Hess}}f(U)[\Omega],\Omega\rangle+\mathcal{O}(t^{3}).$ (61)
The updated unitary matrix is defined by $U(t):=\mathrm{R}_{U}(t\eta)=e^{t\Omega}U$. The corresponding perturbed quantum state, represented by the density operator, is given by
$\psi(t)$ $=U(t)\psi_{0}U(t)^{\dagger}=\left(e^{t\Omega}U\right)\psi_{0}\left(e^{t\Omega}U\right)^{\dagger}$ (62)
$=e^{t\Omega}(U\psi_{0}U^{\dagger})e^{-t\Omega}=e^{t\Omega}\psi e^{-t\Omega}.$ (63)
Expanding this state to the second order in $t$ yields
$\psi(t)$ $=\left(I+t\Omega+\frac{t^{2}}{2}\Omega^{2}\right)\psi\left(I-t\Omega+\frac{t^{2}}{2}\Omega^{2}\right)+\mathcal{O}(t^{3})$
$=\psi+t[\Omega,\psi]+\frac{t^{2}}{2}[\Omega,[\Omega,\psi]]+\mathcal{O}(t^{3}).$ (64)
Substituting $\mathrm{R}_{U}(t\eta)$ into the cost function gives
$f\left(\mathrm{R}_{U}(t\eta)\right)=f(U(t))=\operatorname{Tr}(H\psi(t))$ (65)
$=\operatorname{Tr}(H\psi)+t\operatorname{Tr}(H[\Omega,\psi])+\frac{t^{2}}{2}\operatorname{Tr}(H[\Omega,[\Omega,\psi]])+\mathcal{O}(t^{3}).$ (66)
Note that the zeroth-order term $f(U)=\operatorname{Tr}(H\psi)$. Thus, the remaining task is to verify that the first- and second-order coefficients in Eq.˜61 and Eq.˜66 coincide.
For the first-order coefficient in Eq.˜61, since $\widetilde{\operatorname{grad}}f(U)=[H,\psi]$, we have
$\langle\widetilde{\operatorname{grad}}f(U),\Omega\rangle=-\operatorname{Tr}\big([H,\psi]\Omega\big)=\operatorname{Tr}(H[\Omega,\psi]),$ (67)
where the last equality follows from the identity
$\operatorname{Tr}([A,B]C)=\operatorname{Tr}(A[B,C])$ (68)
for arbitrary $A,B,C$. Therefore, this coefficient coincides with that in Eq.˜66.
For the second-order coefficient in Eq.˜61, we have
$\langle\widetilde{\operatorname{Hess}}f(U)[\Omega],\Omega\rangle=-\operatorname{Tr}(\widetilde{\operatorname{Hess}}f(U)[\Omega]\Omega)$ (69)
$=-\frac{1}{2}\operatorname{Tr}\Big(\underbrace{[H,[\Omega,\psi]]\Omega}_{a:=}+\underbrace{[[H,\Omega],\psi]\Omega}_{b:=}\Big).$ (70)
Using Eq.˜68, the traces of $a$, $b$ can be evaluated as
$\operatorname{Tr}([H,[\Omega,\psi]]\Omega)$ $=\operatorname{Tr}(H[[\Omega,\psi],\Omega])=-\operatorname{Tr}(H[\Omega,[\Omega,\psi]]),$
$\operatorname{Tr}([[H,\Omega],\psi]\Omega)$ $=\operatorname{Tr}([H,\Omega][\psi,\Omega])=-\operatorname{Tr}(H[\Omega,[\Omega,\psi]]).$
Substituting these expressions into Eq.˜70 yields
$\langle\widetilde{\operatorname{Hess}}f(U)[\Omega],\Omega\rangle=\operatorname{Tr}(H[\Omega,[\Omega,\psi]]),$ (71)
which matches the second-order Taylor coefficient in Eq.˜66, thereby verifying the Riemannian Hessian formula of Eq.˜60.
In general manifold optimization, computing a Newton step necessitates the computationally expensive inversion of this Hessian operator. However, as demonstrated in Theorem˜3, when the Hamiltonian is a projector ($H=H^{2}$), the Riemannian gradient strictly emerges as an exact eigenvector of this Hessian operator.
Appendix B Proofs
Proof of Theorem˜3.
Substituting $g=[H,\psi]$ into the definition of $L$ yields
$L(g)=\frac{1}{2}\bigl([H,[g,\psi]]+[[H,g],\psi]\bigr).$ (72)
By Jacobi identity $[A,[B,C]]=[[A,B],C]+[B,[A,C]],$ applied with $A=H$, $B=g$, and $C=\psi$, we obtain
$H,[g,\psi]]=[[H,g],\psi]+[g,g]=[[H,g],\psi].$ (73)
Therefore, $L(g)=[[H,g],\psi].$ Next, since $H^{2}=H$, we compute
$[H,g]$ $=[H,[H,\psi]]=[H,H\psi-\psi H]$ (74)
$=H\psi-2H\psi H+\psi H.$ (75)
Hence, $H,g],\psi]=[H\psi-2H\psi H+\psi H,\psi].$ Using $\psi^{2}=\psi$ and $\psi H\psi=q\psi$, we evaluate the three terms separately:
$[H\psi,\psi]=H\psi-q\psi,\quad[\psi H,\psi]=q\psi-\psi H,$ (76)
$[H\psi H,\psi]=qH\psi-q\psi H.$ (77)
Combining these identities yields
$[[H,g],\psi]$ $=(H\psi-q\psi)-2(qH\psi-q\psi H)+(q\psi-\psi H)$ (78)
$=(1-2q)(H\psi-\psi H)=(1-2q)g.$ (79)
Thus, L(g)=(12q)gL(g)=(1-2q)g. This completes the proof. ∎